# Prompt-aware classifier free guidance for diffusion models

### v1

@@ 1 Introduction
Diffusion model [1] have achieved remarkable success in generative tasks such as image [2], speech [3], and audio [4] synthesis, largely due to their stable training and ability to produce high-fidelity samples. To further improve controllability, CFG [5] has become the dominant technique, where conditional and unconditional predictions are combined to guide the sampling trajectory. By scaling the difference between these two predictions, CFG can greatly enhance alignment with conditioning signals and improve corresponding generation quality in the meantime, leading to substantial improvements in tasks for content generation. Owing to its simplicity and effectiveness, CFG has now become a standard component in diffusion-based generative modeling.
Although diffusion models with proper guidance strategies can already approximate the score of the data distribution and yield strong sampling performance, there remain several opportunities to push their limits, especially for pretrained diffusion backbones. One line of research explores structural inductive biases inside the network to improve sampling without retraining, such as FreeU [6] and MelRefine [7]. Another line focuses on training-free guidance refinements beyond CFG, including Autoguidance [8], TPG [9], and SAG [10], which alleviate the inherent limitations of the unconditional branch in CFG. In addition, a series of works investigate fundamental components of the sampling process itself: Goldnoise [11] examines the role of noise initialization, while CATImage [12] adaptively chooses sampling timesteps based on the type of conditioning signal, aiming for a more optimal trade-off between quality and efficiency. These advances collectively suggest that even for pretrained diffusion models, there is still room to improve controllability, stability, and quality in a training-free manner.
@@ B. Overall Framework
1) Guidance-based dataset generation. For each prompt $p$, multiple generations are produced across candidate scales $\omega\in\mathcal{S}$. Each output is evaluated with modality-specific metrics (KID, CLIP, ImageReward, Precision/Recall for images; AudioBox-Aesthetics for audio). Averaged scores form the oracle label $\mathbf{q}(p,\omega)$.
@@ B. Results and Analysis
Furthermore, Table 3 shows that relying solely on KID and CLIP cannot fully capture overall generation quality and may even yield controversial effects, as improvements in alignment do not necessarily translate to perceptual preference or distributional fidelity. By contrast, combining all four metrics—KID, FID, CLIP, and ImageReward—together with the quadratic regularization term provides a more comprehensive and consistent supervision signal, highlighting the importance of multi-metric integration for stable and reliable scale selection.

### v2

@@ 1 Introduction
Diffusion model [1] have achieved remarkable success in generative tasks such as image [2], speech [3], and audio [4, 5] synthesis, largely due to their stable training and ability to produce high-fidelity samples. To further improve controllability, CFG [6] has become the dominant technique, where conditional and unconditional predictions are combined to guide the sampling trajectory. By scaling the difference between these two predictions, CFG can greatly enhance alignment with conditioning signals and improve corresponding generation quality in the meantime, leading to substantial improvements in tasks for content generation. Owing to its simplicity and effectiveness, CFG has now become a standard component in diffusion-based generative modeling.
Although diffusion models with proper guidance strategies can already approximate the score of the data distribution and yield strong sampling performance, there remain several opportunities to push their limits, especially for pretrained diffusion backbones. One line of research explores structural inductive biases inside the network to improve sampling without retraining, such as FreeU [7] and MelRefine [8]. Another line focuses on training-free guidance refinements beyond CFG, including Autoguidance [9, 10], TPG [11], and SAG [12], which alleviate the inherent limitations of the unconditional branch in CFG. In addition, a series of works investigate fundamental components of the sampling process itself: Goldnoise [13] examines the role of noise initialization, while CATImage [14] adaptively chooses sampling timesteps based on the type of conditioning signal, aiming for a more optimal trade-off between quality and efficiency. These advances collectively suggest that even for pretrained diffusion models, there is still room to improve controllability, stability, and quality in a training-free manner.
@@ B. Overall Framework
1) Guidance-based dataset generation. For each prompt $p$, multiple generations are produced across candidate scales $\omega\in\mathcal{S}$. Each output is evaluated with modality-specific metrics (KID, CLIP, ImageReward [19], Precision/Recall for images; AudioBox-Aesthetics for audio). Averaged scores form the oracle label $\mathbf{q}(p,\omega)$.
@@ B. Results and Analysis
Furthermore, Table 3 shows that relying solely on KID and CLIP cannot fully capture overall generation quality and may even yield controversial effects, as improvements in alignment do not necessarily translate to perceptual preference or distributional fidelity. By contrast, combining all four metrics: KID, FID, CLIP, and ImageReward, together with the quadratic regularization term provides a more comprehensive and consistent supervision signal, highlighting the importance of multi-metric integration for stable and reliable scale selection.